The initial temperature of a body is 80 degree celcius If its temperature falls to 64 degree celcius in 5 min and in 10 min to 52 degree celcius then the temperature of surrounding will be

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

assume newtons law of cooling

dT/dt =-k(T-To)  ...................1

To is the temperature of sorrunding

dT = 64-80 =-16   

so dT/dt  =-16/5  oC /min

mean temperature T = (80 + 64)/2 = 72

so put these value in equation 1

-16/5 = -k(72 -To)   ....................a

for the next 5 sec

mean temperature T = (64 + 52)/2 =58

and dT/dt = (52-64)/5 = -12/5

so put again in equation  1

-12/5  = -k( 58 - To) ............b

now  divide equation a by b

 16/12   = (72-To) /(58-To)

To =16

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin